# Beyond Single-Turn: A Survey on Multi-Turn Interactions with Large Language Models

### v4

@@ 4.5.3 Hyper-Realism and User Perception
An emerging concern with highly advanced dialogue models is fidelity – the interactions can become so human-like that users may struggle to remember (or realize) that they are conversing with an AI. On one hand, human-like fluency and emotional resonance are goals of natural language interfaces; on the other hand, when an AI system crosses a threshold of hyperrealism, it can lead to deception (even if unintentional) and overtrust [273]. Studies have documented users developing strong personal bonds or even a therapeutic alliance with chatbots, with measured alliance levels in some cases approaching those of human therapists [274]. Such emotional engagement can provide comfort or companionship, but it also heightens the danger of overtrust and dependence, potentially displacing human relationships or giving AI undue influence over vulnerable individuals [275, 276]. In response, regulators have emphasized transparency; for example, the proposed EU AI Act mandates that AI systems interacting with people must clearly disclose their artificial identity to prevent deception [276]. Mitigation strategies in design likewise include persistent identity reminders and constrained, less too-human stylistic choices to maintain healthy user boundaries and reduce the illusion of personhood [273].
@@ 5 Conclusion
This survey has provided a comprehensive overview of the rapidly evolving landscape of multi-turn interactions with Large Language Models, making several key contributions to the field. First and foremost, we have introduced a novel task-oriented taxonomy for analyzing multi-turn LLM interactions, departing from the capability-oriented approaches prevalent in existing literature. While previous surveys have focused on isolated capabilities (such as reasoning, memory, or contextual understanding), our framework recognizes that real-world multi-turn applications require the complex collaboration of multiple capabilities working in concert. By categorizing interactions according to tasks—instruction following and conversational engagement—rather than isolated abilities, we offer a more practical taxonomy that better reflects how LLMs are deployed and evaluated in authentic contexts. This task-oriented perspective has enabled us to provide substantial analysis of critical real-world application domains, including role-playing scenarios, healthcare consultations, educational scenario, and LLM jail-breaking where the interplay of multiple capabilities determines performance.
@@
Our analysis demonstrates that multi-turn interactions represent a fundamental paradigm shift in how we utilize and evaluate LLMs. Unlike single-turn interactions, which have dominated early LLM benchmarks, multi-turn settings more closely mirror real-world applications—from sustained dialogues to complex iterative problem-solving. These contexts demand not only factual knowledge but also context retention, coherent reasoning across turns, adaptive behavior, and robust handling of ambiguous or changing user intentions.
@@
The improvement methodologies we’ve extensively reviewed span model-centric approaches (in-context learning, fine-tuning, reinforcement learning, and architectural innovations), external integration strategies (memory augmentation, retrieval mechanisms, and knowledge graphs), and agent-based frameworks (both single-agent and multi-agent systems). By providing this detailed analysis of improvement techniques, we address a significant gap in existing surveys, which have often overlooked the nuanced methodological landscape of multi-turn enhancements.
@@
Despite remarkable progress, significant challenges remain. Context understanding issues persist, with even state-of-the-art models struggling to maintain coherence across extended conversations. Complex reasoning across turns frequently suffers from error propagation and topic-switching difficulties. Adaptation capabilities remain limited, especially regarding dynamic preference learning and knowledge updates during interaction. Our detailed organization of these open challenges provides a roadmap for future research that was previously lacking in the literature.
@@
In conclusion, multi-turn interaction capabilities represent both a frontier challenge and a transformative opportunity for LLM research. By more closely approximating the dynamic, context-dependent nature of human communication, advances in this domain will not only enhance technical performance but also significantly improve the practical utility and trustworthiness of these systems across diverse real-world applications. Our task-oriented taxonomy, extensive examination of real-world domains, detailed analysis of improvement methodologies, and in-depth organization of open challenges collectively provide a foundation for future work in this critical area.

### v5

@@ 4.5.3 Hyper-Realism and User Perception
As dialogue models grow more fluent and emotionally resonant, users may struggle to recognize they are conversing with an AI—raising concerns about unintentional deception and overtrust (Akbulut et al., 2025). Studies document users forming strong personal bonds with chatbots, with therapeutic alliance scores in some cases approaching those of human therapists (Beatty et al., 2022). While such engagement can provide genuine comfort, it also risks fostering dependence and giving AI undue influence over vulnerable individuals (Fang et al., 2025; Boine, 2023). In response, the proposed EU AI Act mandates clear disclosure of artificial identity in human-facing AI systems (Boine, 2023). Design-level mitigations include persistent identity reminders and deliberately less human-like stylistic choices to maintain healthy user boundaries and reduce the illusion of personhood (Akbulut et al., 2025).
@@ 5 Conclusion
This survey has provided a comprehensive overview of the rapidly evolving landscape of multi-turn interactions with Large Language Models, contributing several advances to the field.
@@
We introduced a task-oriented taxonomy for analyzing multi-turn LLM interactions, departing from the capability-oriented frameworks prevalent in existing literature. Prior surveys have largely focused on isolated capabilities—reasoning, memory, or contextual understanding—yet real-world multi-turn applications demand the coordinated interplay of multiple such capabilities. By organizing interactions around tasks—instruction following and conversational engagement—rather than isolated abilities, our taxonomy better reflects how LLMs are deployed and evaluated in practice. This perspective enabled substantive analysis of critical application domains, including role-playing, healthcare consultation, education, and jailbreaking, where performance is determined by the joint operation of many capabilities.
@@
Our analysis establishes that multi-turn interaction represents a fundamental paradigm shift in LLM utilization and evaluation. Unlike single-turn settings that dominated early benchmarks, multi-turn contexts more closely mirror real-world use—from sustained dialogue to complex iterative problem-solving—demanding not only factual knowledge but also context retention, coherent cross-turn reasoning, adaptive behavior, and robust handling of ambiguous or evolving user intent.
@@
The improvement methodologies we surveyed span model-centric approaches (in-context learning, fine-tuning, reinforcement learning, and architectural innovations), external integration strategies (memory augmentation, retrieval mechanisms, and knowledge graphs), and agent-based frameworks (single- and multi-agent systems). This detailed methodological landscape addresses a significant gap in prior surveys, which have often treated multi-turn enhancements only superficially.
@@
Significant challenges nonetheless remain. Even state-of-the-art models struggle to maintain coherence across extended conversations, and complex cross-turn reasoning remains susceptible to error propagation and topic-switching failures. Adaptation capabilities are similarly limited, particularly regarding dynamic preference learning and on-the-fly knowledge updates. Our structured organization of these open challenges provides a concrete roadmap for future research.
@@
Multi-turn interaction capabilities represent both a frontier challenge and a transformative opportunity for LLM research. Advances in this domain will not only improve technical performance but also enhance the practical utility and trustworthiness of LLMs across diverse real-world settings. Our task-oriented taxonomy, domain-level analysis, methodological survey, and organization of open challenges together provide a foundation for future work in this critical area.
Appendix A Appendix
You may include other additional sections here.